# Towards Verifiable Text Generation with Evolving Memory and Self-Reflection

### v2

@@ Appendix B Datasets and Settings
Appendix C Hyper-parameters
@@ Appendix E Prompts
Instruction: Write an accurate, engaging, and concise answer for the given question using only the provided search results (some of which might be irrelevant). Question: {Question} Document: {Document} Answer:
Instructions: You will be provided with a sentence and several related documents. Your task is to directly append citation annotations to the sentence using these documents without changing the sentence. When citing documents, use [1][2][3]. Cite at least one document and at most three documents. If multiple documents support the sentence, only cite a minimum sufficient subset of the documents. Document: {Document} Sentence: {Sentence} Sentence with citation:
Given the original question: {Question}. The context is as follows: {Context}. The claim is: {Claim}. Please generate up to {qg_num} questions that can help verify the claim with the following constraints: 1. You should output no more than {qg_num} questions. 2. The generated questions should be diverse and focus on different aspects of the given claim. Generated questions:
Prompt for Vanilla
Instruction: Write a high-quality answer for the given question using only the provided search results and cite them properly using [1][2][3]. Question: {Question} Document: {Document} Answer:
Prompt for Summ
## Step 1: First Summarize the documents Summarize the following document within 50 words with the question of interest {Question} Return "irrelevant" if the document is “irrelevant" to the question. Try to keep all the important dates, numbers, and names. Title: {Title} Text: {Text} Summary: ## Step 2: Generate the response based on the summary Instruction: Write a high-quality answer for the given question using only the provided search results and cite them properly using [1][2][3]. Question: {Question} Document: {Document} Answer:
Prompt for Snippet
## Step 1: First extract relevant snippet from the documents Given the following passage and the question {Question}, extract a useful span from the passage that can answer the question. Resolve all the coreference issues to make the extracted span understandable and standalone. If the passage is not helpful for answering the question, return “irrelevant". If there are multiple spans, merge them and only output one paragraph. Title: {Title} Text: {Text} Extracted span: ## Step 2: Generate the response based on the snippet Instruction: Write a high-quality answer for the given question using only the provided search results and cite them properly using [1][2][3]. Question: {Question} Document: {Document} Answer:

### v3

@@ Appendix B Datasets and Settings
LLM Evaluation
To ensure a fair comparison of citation quality, we instruct Qwen-MAX to evaluate model generations. Specifically, we assess the quality of the citations in two ways: (1) Citation Recall: The large language model (LLM) is given a sentence and all the passages that the sentence cited, and is asked to judge whether the passages fully support the sentence; (2) Citation Precision: Given a sentence and one of its citations, the LLM is asked to judge whether the citation “fully supports” or “does not support” the sentence. Each citation receives a precision score of 1 if the output sentence has a citation recall of 1 and this citation is “fully support.”
Appendix C Focus Shifting Phenomenon
To analyze the focus shifting phenomenon within the datasets, we employ BGE embedding [^6] https://huggingface.co/BAAI/bge-base-en-v1.5Xiao et al. (2023) to represent each sentence from the LLM’s outputs for all questions. Subsequently, we calculate the cosine similarity between these embedded sentences to construct a similarity matrix for each question. An average of all these similarity matrices is computed, and the results are depicted in Figure 5.
The color intensity on the diagonal of the matrix is the strongest, which signifies a high degree of similarity between each sentence and itself. Moreover, there is a noticeable gradation in color intensity as the distance between the current and the target sentences increases, indicating a decrease in similarity. This pattern illustrates that the content focus of the sentences tends to diverge significantly as the LLM continues the generation, which is also called the focus shifting phenomenonLan and Jiang (2021); Sun et al. (2023a).
(a) Wikihop
(b) WebQ
(c) NQ
(d) ASQA
(e) ELI5
Figure 5: The focus shifting phenomenon.
Appendix D Use of Different NLI models
To further validate the robustness of our method, we conducted additional experiments using two distinct NLI models for inference and evaluation, respectively. Specifically, we employed the t5_11b_trueteacher_and_anli [^7] https://huggingface.co/google/t5_11b_trueteacher_and_anli as the verifier within our VTG system. The t5_xxl_true_nli_mixture [^8] https://huggingface.co/google/t5_xxl_true_nli_mixture was used for evaluation, consistent with the evaluation methodology used in ALCE. We randomly sampled 300 questions from ASQA and NQ datasets and conducted experiments using Vicuna-13B-v1.5-16K as the base LLM. From the results shown in Table 10, we can find that our method can outperform all the baselines in terms of both correctness and citation quality, validating the superiority of our method.
Methods ASQA NQ
Correct.EM Correct.D-F1 Citation.Rec Citation.Prec Citation.F1 Correct.EM Citation.Rec Citation.Prec Citation.F1
Vanilla 30.69 26.95 70.31 59.51 64.46 55.00 70.94 61.20 65.71
Summ 36.26 27.41 57.18 60.44 58.77 56.67 63.28 63.65 63.47
Snippet 35.50 25.92 62.64 57.57 60.00 55.67 55.51 55.71 55.61
Rerank 34.16 26.53 87.75 73.34 79.90 57.00 83.92 72.85 78.00
PostCite 27.80 22.78 34.40 23.30 27.78 43.00 34.02 23.32 27.67
RefineCite 36.04 28.68 73.92 72.50 73.20 54.40 68.40 65.24 66.78
VeriCite 31.89 27.97 77.50 56.50 65.35 55.80 68.37 49.63 57.51
VeriRefine 30.04 26.40 75.67 64.83 69.83 55.20 74.93 61.21 67.38
Vtg 40.28 29.22 92.97 88.00 90.42 59.00 87.24 80.66 83.82
Table 10: Performance comparison across ASQA and NQ datasets
Appendix E Hyper-parameters
@@ Appendix G Prompts
Prompt for Sentence Generator Instruction: Write an accurate, engaging, and concise answer for the given question using only the provided search results (some of which might be irrelevant). Question: {Question} Document: {Document} Answer:
Prompt for Citation Generator Instructions: You will be provided with a sentence and several related documents. Your task is to directly append citation annotations to the sentence using these documents without changing the sentence. When citing documents, use [1][2][3]. Cite at least one document and at most three documents. If multiple documents support the sentence, only cite a minimum sufficient subset of the documents. Document: {Document} Sentence: {Sentence} Sentence with citation:
Prompt for Query Generation Given the original question: {Question}. The context is as follows: {Context}. The claim is: {Claim}. Please generate up to {qg_num} questions that can help verify the claim with the following constraints: 1. You should output no more than {qg_num} questions. 2. The generated questions should be diverse and focus on different aspects of the given claim. Generated questions:
Prompt for Vanilla Instruction: Write a high-quality answer for the given question using only the provided search results and cite them properly using [1][2][3]. Question: {Question} Document: {Document} Answer:
Prompt for Summ ## Step 1: First Summarize the documents Summarize the following document within 50 words with the question of interest {Question} Return "irrelevant" if the document is “irrelevant" to the question. Try to keep all the important dates, numbers, and names. Title: {Title} Text: {Text} Summary: ## Step 2: Generate the response based on the summary Instruction: Write a high-quality answer for the given question using only the provided search results and cite them properly using [1][2][3]. Question: {Question} Document: {Document} Answer:
Prompt for Snippet ## Step 1: First extract relevant snippet from the documents Given the following passage and the question {Question}, extract a useful span from the passage that can answer the question. Resolve all the coreference issues to make the extracted span understandable and standalone. If the passage is not helpful for answering the question, return “irrelevant". If there are multiple spans, merge them and only output one paragraph. Title: {Title} Text: {Text} Extracted span: ## Step 2: Generate the response based on the snippet Instruction: Write a high-quality answer for the given question using only the provided search results and cite them properly using [1][2][3]. Question: {Question} Document: {Document} Answer:
Prompt for Citation Evaluation **Role: Data Annotator** **Instructions:** You are provided with the following materials: - **Passage**: passage - **Sentence**: sentence **Task**: Assess whether the passage fully supports the sentence. **Choices**: 1. **Fully Supports**: Select this option if the passage completely and clearly supports every aspect of the sentence. 2. **Does Not Fully Support**: Select this option if any discrepancies, omissions, or inaccuracies in the passage prevent it from fully supporting the sentence. **Output**: - If the passage fully supports the sentence, output "Yes." - If it does not, output "No." **Note**: Please refrain from adding any content not requested in the instructions.